Jesse has 38 nickels and dimes. The number of nickels is 1 less than twice the number of dimes. How many of each coin does Jesse have?
Answer by fractalier(6550)

Call the numbers of nickels and dimes Jesse has, n and d. Then we can write
n + d = 38 and
n = 2d - 1
Substitute that into the first equation and get
2d-1 + d = 38
3d - 1 = 38
3d = 39
d = 13 dimes so that
n = 2(13) - 1 = 25 nickels